Can you find an example that the Gauss-Seidel iteration converges and the Jacobi iteration diverges?

Hint: P(A) of Jacobi >1, but P(A) of Gauss-Seidel <1, or P(A) of Gauss >1, but P(A) of jacobi <1.
Answer #1

3,-5X2 = -4 G1,92) = (0,0). 721-22 = 6 213-07522 72. = -6 + 7211 in 0 1 2 3 4 2 0 -4 -34 -174 -1244 22 0 -6 -34 -244 -1244 el